OpenStudy (anonymous):

If u=10i+9j and v=91-4j find u.v

OpenStudy (swissgirl):

I think its dot product

OpenStudy (anonymous):

I am pretty sure that he is referring to vectors

OpenStudy (jhannybean):

oh LOL my bad.....

OpenStudy (anonymous):

@csmith0355 The dot-product of two vectors \(\bf \vec u=(u_{1},u_{2})\) and \(\bf \vec v=(v_{1},v_{2})\) is given by:\[\bf \vec u \frac{}{} \vec v=u_{1}v_{1}+u_{2}v_{2}=|\vec u||\vec v|\cos(\theta)\], where theta is the angle between the vectors.

OpenStudy (jhannybean):

Yeah,.... i'm not familiar with this stuff yet :(

OpenStudy (jhannybean):

Sorry if i confused you.

OpenStudy (anonymous):

To start you off, I will reveal the following information/hints to you:\[\bf u=10i+9j=[10,0]+[0,9]=[10,9]\]@csmith0355 Is the second vector v = 91 - 4j?

OpenStudy (anonymous):

@csmith0355

OpenStudy (anonymous):

i got 54

OpenStudy (anonymous):

and yes this is vectors sorry 4 the late response